# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
IntroductionElectron-phonon coupling (EPC) is fundamental to understanding a range of phenomena in condensed matter physics and materials science Giustino (2017); Grimvall (1981), including phonon-mediated superconductivity Pellegrini and Sanna (2024), electrical resistivity in metals, carrier lifetime and mobility in semiconductors, and the temperature-dependent behavior of electronic structures Engel et al. (2020, 2022). The development of Kohn-Sham (KS) density functional theory (DFT) has enabled the quantitative prediction of EPCs through first-principles calculations, which have advanced significantly over the last few decades Giustino (2017). The two main methods exist for calculating EPC matrix elements are density functional perturbation theory (DFPT) Baroni et al. (2001) and the finite-difference method Shang and Yang (2023); Monserrat (2018).
@@
DFPT, often combined with maximally localized Wannier functions Giustino et al. (2007); Zhou et al. (2021a), provides an efficient means of calculating EPC matrix elements across the entire Brillouin Zone (BZ). However, it is typically limited to traditional density functional approximations (DFAs) that are not explicitly dependent on KS orbitals, such as the local-density approximation (LDA) and generalized-gradient approximations (GGAs). The finite-difference approach offers greater flexibility, allowing the inclusion of orbital-dependent density functionals and integration with advanced electronic structure solvers that consider higher-order interactions  Monserrat (2018); Antonius et al. (2014); Yin et al. (2013), but it often requires large supercells, making it computationally intensive.
@@
For complex materials with $d$- and $f$-electrons, LDA and GGAs often fail to accurately capture their electronic structures Furness et al. (2018); Lane et al. (2018); Zhang et al. (2020a, 2024) and phonon properties Ning et al. (2024, 2022, 2023); Zhou et al. (2021b), resulting in limited predictions for EPC. While methods like DFPT+$U$ can improve accuracy Zhou et al. (2021b); Floris et al. (2020), they rely on parameterizations specific for $d$- and $f$-electrons. More precise approaches, such as combining GW with DFPT Li et al. (2019, 2021), further enhance EPC predictions but involve substantial computational costs. This underscores the need for efficient methods that can deliver accurate EPC predictions without the computational burden of beyond-DFT methods.
@@
Meta-GGAs like the strongly-constrained and appropriately-normed (SCAN)  Sun et al. (2015) density functional and its variant r2SCAN Furness et al. (2020), which are orbital dependent and scale the same computationally as LDA and GGAs, have shown promise in providing accurate descriptions of total energy and electronic structures across diverse systems Sun et al. (2015); Furness et al. (2020); Sun et al. (2016). These functionals have proven effective in materials ranging from liquid water Chen et al. (2017) to metal oxides Gautam and Carter (2018), and complex materials with $d$- and $f$-electrons Zhang et al. (2020a, 2022), like the cuprates Lane et al. (2018); Furness et al. (2018); Zhang et al. (2020b) and nickelates Zhang et al. (2021); Lane et al. (2023); Zhang et al. (2024). Recent work also indicates that SCAN/r2SCAN can accurately calculate phonon dispersions of cuprates Zhang et al. (2020a); Ning et al. (2023) and NiO Ning et al. (2022). However, their performance on EPC predictions remains unexplored.
@@
Here, we apply the r2SCAN meta-GGA within the finite-difference approach Engel et al. (2020, 2022) to calculate EPC of transition metal compounds including cobalt oxide (CoO) and nickel oxide (NiO), as well as a main-group conventional superconductor, magnesium diboride (MgB2). We show that r2SCAN correctly captures strong EPC effects in CoO and NiO, significantly improving over the standard Perdew-Burke-Ernzerhof (PBE) GGA without requiring the introduction of empirical parameters. At the same time, r2SCAN achieves high accuracy in modeling phonon-mediated superconducting properties of MgB2.
@@
Evaluation of EPCs in CoO and NiOCoO and NiO are two fundamental transition metal oxides that pose challenges for DFT Zhang et al. (2020a); Trimarchi et al. (2018). In our simulations, we use the low-temperature antiferromagnetic (AFM) phase of CoO and NiO, characterized by ferromagnetic planes along the [111] direction that alternate in magnetization. Upon relaxation, the cubic rock-salt structures of CoO and NiO exhibit slight distortions due to the AFM order. Experimentally, CoO and NiO are polar semiconductors, which exhibit long-range dipole-dipole interactions arising from the polarization of collective ionic movements. These interactions can cause splitting between longitudinal optical (LO) and transverse optical (TO) phonon modes at the $\Gamma$ point—a phenomenon known as LO-TO splitting. To account for this effect, the nonanalytic contributions needed for calculating phonon dispersions and EPCs must be included Baroni et al. (2001); Giannozzi et al. (1991), which depend on the Born effective charges ($Z^{*}$) and the ion-clamped macroscopic dielectric tensor ($\epsilon^{\infty}$).
@@
Figure 1(a) presents the phonon dispersions of CoO calculated using r2SCAN and PBE, along with the EPC matrix elements from r2SCAN. PBE predicts unphysical negative phonon frequencies along the $M$-$F$-$\Gamma$ and $T$-$K$-$L$ paths, indicating lattice instability in CoO, as reported in previous studies Floris et al. (2020); Zhou et al. (2021b). Consequently, PBE is unsuitable for calculating EPC in CoO. The PBE+$U$ method is often applied to eliminate these negative frequencies, allowing DFPT+$U$ to provide accurate EPC predictions Floris et al. (2020); Zhou et al. (2021b); Wdowik and Parlinski (2007). Remarkably, r2SCAN resolves these lattice instabilities without requiring any $U$ parameters, as shown in Fig. 1(a). Furthermore, a comparison of phonon dispersions calculated by r2SCAN and PBE reveals that r2SCAN shifts phonon frequencies—particularly for the optical modes—upward, achieving better agreement with experimental results, as shown in Figs. 1(a) and S2 of the Supplementary Materials (SM).
@@
In Fig. 1(a), our r2SCAN calculations also reveal significant LO-TO splitting in CoO. The predicted values of $\omega_{\text{LO}}=(59.4,\,62.6)$ meV and $\omega_{\text{TO}}=(38.5,\,45.7)$ meV from r2SCAN are close to experimental measurements [see Fig. S2]. Additionally, we observe notable splitting between the TO1 and TO2 modes along the $F-\Gamma-T$ direction [Fig. 1(a)]. This splitting primarily results from the structural distortion induced by AFM order, which breaks the equivalence along the [111] directions ($\Gamma-T$), as noted in previous studies Floris et al. (2020, 2011). A similar, though smaller, LO mode splitting is observed at the $\Gamma$ point. Similar results were obtained using PBE+$U$ calculations Floris et al. (2020), although the top two optical modes along the $F-\Gamma-T$ direction are reversed in our r2SCAN results compared to PBE+$U$, likely due to the larger Born effective charge $Z^{*}$ (2.25) and ion-clamped macroscopic dielectric function $\epsilon^{\infty}$ (6.17) values predicted by PBE+$U$ Floris et al. (2020).
@@
Table 1: The calculated band gap ($E_{g}$), diagonals of the Born effective charge ($Z^{*}$), local magnetic moments (m) and ion-clamped macroscopic dielectric constants in the independent-particle approximation ($\epsilon^{\infty}_{\text{IP}}$) and including $f_{xc}$ ($\epsilon^{\infty}$) of CoO and NiO from PBE and r2SCAN functionals, as well as measured experimental values.
Figure 1(a) also shows that our r2SCAN calculations reveal substantial EPC contributions from both LO and TO modes near the $\Gamma$ point, with maximum EPC matrix elements [g defined in Eq. (1)] exceeding $10^{4}$ meV. This result aligns with previous DFPT+$U$ findings Zhou et al. (2021b) quantitatively. In transition metal oxides like CoO and NiO, the polar nature of the bonds leads to strong interactions between electrons and LO and TO phonons via the Fröhlich mechanism Fröhlich (1954). These interactions become particularly significant as the wave vector $\mathbf{q}\rightarrow 0$, causing a divergence in the EPC matrix elements for the LO and TO modes [Fig. 1 (a)]. These interactions are often substantial enough to form large polarons—quasiparticles composed of an electron or hole coupled with lattice distortions—which play a crucial role in charge transport and the dynamic behavior of electrons in polar materials. It is therefore significant for r2SCAN to capture the Fröhlich interaction in polar materials without relying on empirical parameters Zhou et al. (2021b, 2018); Zhou and Bernardi (2016).
The phonon dispersions of NiO calculated using r2SCAN and PBE are shown in Fig. 1(b), where a similar phonon softening trend from PBE is observed, as seen in CoO. Additionally, r2SCAN accurately captures the LO-TO mode splitting in NiO due to the inclusion of the nonanalytic term. Similar to CoO, r2SCAN calculations reveal a clear splitting of the TO1 and TO2 modes at the $\Gamma$ point in NiO, though the LO mode splitting is comparatively smaller, as discussed in Ref. Floris et al. (2011). Notably, the TO mode splitting in NiO is smaller than in CoO [Fig.1(b)], which can be attributed to the smaller local magnetic moments of NiO [Table 1], as pointed out in the previous study Floris et al. (2020). Importantly, the phonon characteristics computed with r2SCAN show strong consistency with those obtained using PBE+$U$, as reported in Ref. Floris et al. (2011).
@@
Why does r2SCAN yield superior EPC predictions in CoO and NiO? The results above clearly demonstrate that r2SCAN significantly improves the accuracy of EPC descriptions for CoO and NiO over PBE. Our analysis indicates that the superior performance of r2SCAN over PBE can largely be attributed to the more exact constraints that r2SCAN satisfies Sun et al. (2015, 2016); Furness et al. (2020). Additionally, for transition and rare-earth metal compounds, r2SCAN exhibits reduced self-interaction error (SIE) compared to PBE, significantly improving the description of their electronic structures [Fig. 2 and S1] Zhang et al. (2020b, 2022). The significant improvement of r2SCAN over PBE in describing the phonon characteristics of NiO, particularly for the optical modes, has also been attributed to the reduction of SIE Ning et al. (2022). SIE arises from the imperfect cancellation of the spurious classical Coulomb self-interaction by the approximate exchange-correlation self-interaction Perdew and Zunger (1981). Since the repulsive self-Coulomb term exceeds the attractive self-exchange-correlation term, the net SIE is generally positive, causing orbitals to be underbound (with orbital energies too high) and wave functions to be excessively delocalized. This leads to electron densities that are overly spread out and easily perturbed, ultimately resulting in underestimated magnetic moments and band gaps. The reduction of SIE in r2SCAN compared to PBE therefore results in less delocalization error, more compact $d$- and $f$-orbitals, and fewer fractional occupations, which stabilize the magnetic moments Zhang et al. (2020a, 2022) and increase band gaps for NiO and CoO as shown in Table 1.
@@
Table 1 also presents the calculated $Z^{*}$ and $\epsilon^{\infty}$ values for CoO and NiO from both PBE and r2SCAN, compared with experimental data. Both $Z^{*}$ and $\epsilon^{\infty}$ are critical quantities used in the nonanalytic contributions needed for calculating phonon dispersions and EPCs. The Born effective charge $Z^{*}$ measures the dynamic charge of an atom in a crystal, capturing how polarization changes in response to atomic displacement and accounting for the redistribution of electronic charge under an applied electric field Ghosez et al. (1998); Gonze and Lee (1997); King-Smith and Vanderbilt (1993); Resta (1993). The ion-clamped macroscopic dielectric tensor $\epsilon^{\infty}$ describes the dielectric response of a material, considering only the electronic contribution without ionic motion, reflecting how the electronic cloud in a crystal responds to an external electric field Ghosez et al. (1998); Gonze and Lee (1997); King-Smith and Vanderbilt (1993); Resta (1993). $Z^{*}$ and $\epsilon^{\infty}$ are thus expected to be overestimated by SIE that leads to overly delocalized and easily perturbed electron densities. This can be seen by comparing the $Z^{*}$ and $\epsilon^{\infty}$ values calculated from r2SCAN and PBE against the experimental data. For CoO, we find that the r2SCAN predictions for $Z^{*}$ and $\epsilon^{\infty}$ are in good agreement with experimental results Sakurai et al. (1968); Gielisse et al. (1965). However, PBE incorrectly predicts a metallic ground state for CoO as shown in Fig. S1(b) to be discussed later, resulting in the vanishing of the nonanalytic terms. For NiO, PBE significantly overestimates $\epsilon^{\infty}$ by a factor of 3.77, whereas r2SCAN provides significant improvement, reducing the overestimation to 31.93%. The calculated $Z^{*}$ value using r2SCAN is 2.23, which is in excellent agreement with the experimental value of 2.2, while PBE overestimates $Z^{*}$ by 8.18%. These results corroborate that r2SCAN reduces SIE in comparison with PBE for NiO and CoO.
@@
The above analysis strongly indicates that the improvement of r2SCAN over PBE on EPC matrix elements originates from the SIE reduction too. In the context of the Kohn-Sham (KS) DFT Giustino (2017), the EPC matrix element is defined as:
@@
where $v^{\text{en}}$ is the potential energy arising from electron-nuclei interactions, $v^{\text{H}}$ is the Hartree potential, and $v^{\text{xc}}$ is the exchange-correlation potential. This can be recast into
$\partial_{\mathbf{q}\nu}v^{\text{KS}}=(\epsilon^{\text{Hxc}})^{-1}\partial_{\mathbf{q}\nu}v^{\text{en}}.$ (3)
@@
$\epsilon^{\text{Hxc}}=1-(v^{\text{C}}+f^{\text{xc}})\chi^{0}.$ (4)
@@
Figure 2: The calculated electronic structure of $G$-type AFM phase of CoO and NiO using r2SCAN functional. (a) Orbitally-projected band structure of CoO along high-symmetry lines in the BZ, where the Co $t_{2g}$ and $e_{g}$, and O $p$ orbitals are marked with blue, red, and green colors, respectively. The side panel shows the partial density of states of the respective orbitals. (b) Same as (a), but for NiO.
The ion-clamped macroscopic dielectric constants $\epsilon^{\infty}$ are quantities related to $\epsilon^{\text{Hxc}}$. For a given density functional, $\epsilon^{\infty}$ can be calculated in the independent particle picture where $f^{\text{xc}}$ is ignored, with the value denoted as $\epsilon^{\infty}_{\text{IP}}$. Table 1 shows that the difference between $\epsilon^{\infty}_{\text{IP}}$ and $\epsilon^{\infty}$ of a density functional is much smaller than the difference between $\epsilon^{\infty}$ of different density functionals, suggesting that $f^{\text{xc}}$ has a less important effect on $\epsilon^{\infty}$ than the $\chi_{0}$ effect of the chosen density functional. We note that r2SCAN satisfies more exact constraints than PBE does with more accurate total energies in general, suggesting that r2SCAN can have a better $f^{\text{xc}}$ Nazarov and Vignale (2011).
The KS linear response function $\chi_{0}$ describes how the electron density $n$ reacts to small changes in the Kohn-Sham potential $v^{\text{KS}}$. This linear response function is often overestimated in magnitude by SIE, which leads to an overly delocalized and easily perturbed electron density, as well as an underestimated band gap that is inversely proportional to the linear response function.
For CoO, PBE fails to open a band gap, as illustrated in Fig. S1(b). This leads to too large $\chi_{0}$, resulting in dynamical instabilities with the appearance of imaginary frequencies in the phonon calculations as shown in Figure 1(a). In contrast, due to the SIE reduction, r2SCAN predicts an indirect band gap of 0.85 eV, attributed to the splitting between the occupied $t_{2g}$ orbitals and the unoccupied $t_{2g}$ orbitals in the spin-down channel, as shown in Fig. 2(a). This undoubtedly improves $\chi_{0}$ and thus provides an improved phonon dispersion as well as strong EPC effects at the LO and TO phonon modes around the $\Gamma$ point as shown in Figure 1(a). Since a smaller gap generally correlates with a larger $\epsilon^{\infty}$ Penn (1962); Onishi and Fu (2024), the improved gaps from r2SCAN explain our values of $\epsilon^{\infty}$ being closer to experimental results as shown in Table. 1.

### v2

@@ Abstract
IntroductionElectron-phonon coupling (EPC) is fundamental to understanding a range of phenomena in condensed matter physics and materials science Giustino (2017); Grimvall (1981), including phonon-mediated superconductivity Pellegrini and Sanna (2024), electrical resistivity in metals, carrier lifetime and mobility in semiconductors, and the temperature-dependent behavior of electronic structures Engel et al. (2020, 2022). The development of Kohn-Sham (KS) density functional theory (DFT) has enabled the quantitative prediction of EPCs through first-principles calculations, which have advanced significantly over the last few decades Giustino (2017). The two main methods that exist for calculating EPC matrix elements are density functional perturbation theory (DFPT) Baroni et al. (2001) and the finite-difference method Shang and Yang (2023); Monserrat (2018).
@@
DFPT, often combined with maximally localized Wannier functions Giustino et al. (2007); Zhou et al. (2021a), provides an efficient means of calculating EPC matrix elements across the entire Brillouin Zone (BZ). However, it is typically limited to traditional density functional approximations (DFAs) that are not explicitly dependent on KS orbitals, such as the local-density approximation (LDA) and generalized-gradient approximations (GGAs). The finite-difference approach offers greater flexibility, allowing the inclusion of orbital-dependent density functionals and integration with advanced electronic structure solvers that consider correlation effects beyond the mean-field description Monserrat (2018); Antonius et al. (2014); Yin et al. (2013), but it often requires large supercells, making it computationally intensive.
@@
For complex materials with $d$- and $f$-electrons, LDA and GGAs often fail to accurately capture their electronic structures Furness et al. (2018); Lane et al. (2018); Zhang et al. (2020a, 2024) and phonon properties Ning et al. (2024, 2022, 2023); Zhou et al. (2021b), resulting in inaccurate predictions for EPCs. While methods like DFPT+$U$ can improve accuracy Zhou et al. (2021b); Floris et al. (2020), they rely on parameterizations specific for $d$- and $f$-electrons. More precise approaches, such as combining GW with DFPT Li et al. (2019, 2021), further enhance EPC predictions but involve substantial computational costs. This underscores the need for efficient methods that can deliver accurate EPC predictions without the computational burden of beyond-DFT methods.
@@
Meta-GGAs like the strongly-constrained and appropriately-normed (SCAN)  Sun et al. (2015) density functional and its variant r2SCAN Furness et al. (2020), which are orbital dependent and scale computationally similar as LDA and GGAs, have shown promise in providing accurate descriptions of total energies and electronic structures across diverse systems Sun et al. (2015); Furness et al. (2020); Sun et al. (2016). These functionals have proven effective in materials ranging from liquid water Chen et al. (2017) to metal oxides Gautam and Carter (2018), and complex materials with $d$- and $f$-electrons Zhang et al. (2020a, 2022), like the cuprates Lane et al. (2018); Furness et al. (2018); Zhang et al. (2020b) and nickelates Zhang et al. (2021); Lane et al. (2023); Zhang et al. (2024). Recent work also indicates that SCAN/r2SCAN can accurately calculate phonon dispersions of cuprates Zhang et al. (2020a); Ning et al. (2023) and NiO Ning et al. (2022). However, their performance on EPC predictions remains unexplored.
@@
Here, we apply the r2SCAN meta-GGA within the finite-difference approach Engel et al. (2020, 2022) to calculate EPCs of transition metal compounds including cobalt oxide (CoO) and nickel oxide (NiO), as well as a main-group conventional superconductor, magnesium diboride (MgB2). We show that r2SCAN correctly captures strong EPC effects in CoO and NiO, significantly improving over the standard Perdew-Burke-Ernzerhof (PBE) GGA without requiring the introduction of empirical parameters. At the same time, r2SCAN achieves high accuracy in modeling phonon-mediated superconducting properties of MgB2.
@@
Evaluation of EPCs in CoO and NiOCoO and NiO are two fundamental transition metal oxides that pose challenges for DFT Zhang et al. (2020a); Trimarchi et al. (2018). In our simulations, we use the low-temperature antiferromagnetic (AFM) phase of CoO and NiO, characterized by ferromagnetic planes along the [111] direction that alternate in magnetization. Upon relaxation, the cubic rock-salt structures of CoO and NiO exhibit slight distortions due to the AFM order. Experimentally, CoO and NiO are polar semiconductors, which exhibit long-range dipole-dipole interactions arising from the polarization of collective ionic movements. These interactions cause splittings between longitudinal optical (LO) and transverse optical (TO) phonon modes at the $\Gamma$ point—a phenomenon known as LO-TO splitting. To account for this effect, the nonanalytic contributions needed for calculating phonon dispersions and EPCs must be included Baroni et al. (2001); Giannozzi et al. (1991), which depend on the Born effective charges ($Z^{*}$) and the ion-clamped macroscopic dielectric tensor ($\epsilon^{\infty}$).
@@
Figure 1(a) presents the phonon dispersions of CoO calculated using r2SCAN and PBE, along with the EPC matrix elements from r2SCAN. PBE predicts unphysical negative phonon frequencies along the $M$-$F$-$\Gamma$ and $T$-$K$-$L$ paths, indicating lattice instability in CoO, as reported in previous studies Floris et al. (2020); Zhou et al. (2021b). Consequently, PBE is unsuitable for calculating the EPC in CoO. The PBE+$U$ method is often applied to eliminate these negative frequencies, allowing DFPT+$U$ to provide accurate EPC predictions Floris et al. (2020); Zhou et al. (2021b); Wdowik and Parlinski (2007). Remarkably, r2SCAN resolves these lattice instabilities without requiring any $U$ parameters, as shown in Fig. 1(a). Furthermore, a comparison of phonon dispersions calculated by r2SCAN and PBE reveals that r2SCAN shifts phonon frequencies—particularly for the optical modes—upward, achieving better agreement with experimental results, as shown in Figs. 1(a) and S2 of the Supplementary Materials (SM).
@@
In Fig. 1(a), our r2SCAN calculations also reveal significant LO-TO splitting in CoO. Note that PBE doesn’t capture this LO-TO splitting because it incorrectly predicts a metallic ground state for CoO [Fig. S1 (b)]. The predicted values of $\omega_{\text{LO}}=(59.4,\,62.6)$ meV and $\omega_{\text{TO}}=(38.5,\,45.7)$ meV from r2SCAN are close to experimental measurements [see Fig. S2]. Additionally, we observe notable splitting between the TO1 and TO2 modes along the $F-\Gamma-T$ direction [Fig. 1(a)]. This splitting primarily results from the structural distortion induced by the AFM order, which breaks the equivalence along the [111] directions ($\Gamma-T$), as noted in previous studies Floris et al. (2020, 2011). A similar, though smaller, LO mode splitting is observed at the $\Gamma$ point. Similar results were obtained using PBE+$U$ calculations Floris et al. (2020), although the top two optical modes along the $F-\Gamma-T$ direction are reversed in our r2SCAN results compared to PBE+$U$, likely due to the larger Born effective charge $Z^{*}$ (2.25) and ion-clamped macroscopic dielectric function $\varepsilon^{\infty}$ (6.17) values predicted by PBE+$U$ Floris et al. (2020).
@@
Table 1: The calculated band gap ($E_{g}$), diagonals of the Born effective charge tensor ($Z^{*}$), local magnetic moments (m) and ion-clamped macroscopic dielectric constants in the RPA ($\varepsilon^{\infty}_{\text{RPA}}$) and including $f_{xc}$ ($\varepsilon^{\infty}$) of CoO and NiO for the PBE and r2SCAN functionals, as well as measured experimental values.
Figure 1(a) also shows that our r2SCAN calculations reveal substantial EPC contributions from both LO and TO modes near the $\Gamma$ point, with maximum EPC matrix elements [g defined in Eq. (1)] exceeding $10^{4}$ meV. This result aligns with previous DFPT+$U$ findings Zhou et al. (2021b) quantitatively. In transition metal oxides like CoO and NiO, the polar nature of the bonds leads to strong interactions between electrons and LO and TO phonons via the Fröhlich mechanism Fröhlich (1954). These interactions become particularly significant as the wave vector $\mathbf{q}\rightarrow 0$, causing a divergence in the EPC matrix elements for the LO and TO modes [Fig. 1 (a)]. These interactions are often substantial enough to form large polarons—quasiparticles composed of an electron or hole coupled with lattice distortions—these play a crucial role in charge transport and the dynamic behavior of electrons in polar materials. It is a significant advantage of r2SCAN to capture the Fröhlich interaction in polar materials without relying on empirical parameters Zhou et al. (2021b, 2018); Zhou and Bernardi (2016).
The phonon dispersions of NiO calculated using r2SCAN and PBE are shown in Fig. 1(b), where a similar phonon softening trend from r2SCAN to PBE is observed, as seen in CoO. Additionally, r2SCAN accurately captures the LO-TO mode splitting in NiO due to the inclusion of the nonanalytic term. Similar to CoO, r2SCAN calculations reveal a clear splitting of the TO1 and TO2 modes at the $\Gamma$ point in NiO, though the LO mode splitting is comparatively smaller, as discussed in Ref. Floris et al. (2011). Notably, the TO mode splitting in NiO is smaller than in CoO [Fig.1(b)], which can be attributed to the smaller local magnetic moments of NiO [Table 1], as pointed out in the previous study Floris et al. (2020). Importantly, the phonon characteristics computed with r2SCAN show strong consistency with those obtained using PBE+$U$, as reported in Ref. Floris et al. (2011).
@@
Why does r2SCAN yield superior EPC predictions in CoO and NiO? The results above clearly demonstrate that r2SCAN significantly improves the accuracy of EPC descriptions for CoO and NiO over PBE. Our analysis indicates that the superior performance of r2SCAN over PBE can largely be attributed to the more exact constraints that r2SCAN satisfies Sun et al. (2015, 2016); Furness et al. (2020). Additionally, for transition and rare-earth metal compounds, r2SCAN exhibits reduced a self-interaction error (SIE) compared to PBE, significantly improving the description of their electronic structures [Fig. 2 and S1] Zhang et al. (2020b, 2022). The significant improvement of r2SCAN over PBE in describing the phonon characteristics of NiO, particularly for the optical modes, has also been attributed to the reduction of the SIE Ning et al. (2022). SIE arises from the imperfect cancellation of the spurious classical Coulomb self-interaction by the approximate exchange-correlation self-interaction Perdew and Zunger (1981). Since the repulsive self-Coulomb term exceeds the attractive self-exchange-correlation term, the net SIE is generally positive for semi-local functionals, causing orbitals to be underbound (with orbital energies too high) and wave functions to be excessively delocalized. This leads to electron densities that are overly spread out and easily perturbed, ultimately resulting in underestimated magnetic moments and band gaps. The reduction of the SIE in r2SCAN compared to PBE therefore results in less delocalization error, more compact $d$- and $f$-orbitals, and fewer fractional occupations, which stabilize the magnetic moments Zhang et al. (2020a, 2022) and increase the band gaps for NiO and CoO as shown in Table 1.
@@
Table 1 also presents the calculated $Z^{*}$ and $\varepsilon^{\infty}$ values for CoO and NiO from both PBE and r2SCAN, compared with experimental data. Both $Z^{*}$ and $\varepsilon^{\infty}$ are critical quantities used in the nonanalytic contributions needed for calculating phonon dispersions and EPCs. The Born effective charge $Z^{*}$ measures the dynamic charge of an atom in a crystal, capturing how polarization changes in response to the atomic displacement and accounting for the redistribution of electronic charge under an applied electric field Ghosez et al. (1998); Gonze and Lee (1997); King-Smith and Vanderbilt (1993); Resta (1993). The ion-clamped macroscopic dielectric tensor $\varepsilon^{\infty}$ describes the long-range dielectric response of a material, considering only the electronic contribution without ionic motion, reflecting how the electronic cloud in a crystal responds to an external electric field Ghosez et al. (1998); Gonze and Lee (1997); King-Smith and Vanderbilt (1993); Resta (1993). $Z^{*}$ and $\varepsilon^{\infty}$ are expected to be overestimated due to the SIE that leads to overly delocalized and easily perturbed electron densities. This can be seen by comparing the $Z^{*}$ and $\varepsilon^{\infty}$ values calculated from r2SCAN and PBE against the experimental data. For CoO, we find that the r2SCAN predictions for $Z^{*}$ and $\varepsilon^{\infty}$ are in good agreement with experimental results Sakurai et al. (1968); Gielisse et al. (1965). However, PBE incorrectly predicts a metallic ground state for CoO as shown in Fig. S1(b), to be discussed later, resulting in the vanishing of the nonanalytic terms. For NiO, PBE significantly overestimates $\varepsilon^{\infty}$ by a factor of 3.77, whereas r2SCAN provides significant improvement, reducing the overestimation to 31.93%. The calculated $Z^{*}$ value using r2SCAN is 2.23, which is in excellent agreement with the experimental value of 2.2, while PBE overestimates $Z^{*}$ by 8.18%. These results corroborate that r2SCAN reduces SIE in comparison with PBE for NiO and CoO.
@@
The above analysis strongly indicates that the improvement of r2SCAN over PBE on EPC matrix elements originates from the SIE reduction. In the context of the Kohn-Sham (KS) DFT Giustino (2017), the EPC matrix element is defined as:
@@
where $v^{\text{en}}$ is the potential arising from electron-nuclei interactions, $v^{\text{H}}$ is the Hartree potential, and $v^{\text{xc}}$ is the exchange-correlation potential. This can be recast into Onida et al. (2002):
$\partial_{\mathbf{q}\nu}v^{\text{KS}}=(\varepsilon^{\text{Hxc}})^{-1}\partial_{\mathbf{q}\nu}v^{\text{en}}.$ (3)
@@
$\varepsilon^{\text{Hxc}}=1-(v^{\text{C}}+f^{\text{xc}})\chi^{0}.$ (4)
@@
Figure 2: The calculated electronic structure of $G$-type AFM phase of CoO and NiO using the r2SCAN functional. (a) Orbitally-projected band structure of CoO along high-symmetry lines in the BZ, where the Co $t_{2g}$ and $e_{g}$, and O $p$ orbitals are marked with blue, red, and green colors, respectively. The side panel shows the partial density of states of the respective orbitals. (b) Same as (a), but for NiO.
The ion-clamped macroscopic dielectric constants $\varepsilon^{\infty}$ are quantities related to the head ($\mathbf{q}\rightarrow 0$) of $\varepsilon^{\text{Hxc}}$ Gajdoš et al. (2006). For a given density functional, $\varepsilon^{\infty}$ can be calculated using the Random Phase Approximation (RPA) where $f^{\text{xc}}$ is ignored, with the value denoted as $\varepsilon^{\infty}_{\text{RPA}}$ Gajdoš et al. (2006); Paier et al. (2008). Table 1 shows that the difference between $\varepsilon^{\infty}_{\text{RPA}}$ and $\varepsilon^{\infty}$ of a density functional is much smaller than the difference between $\varepsilon^{\infty}$ of different density functionals, suggesting that $f^{\text{xc}}$ has a less important effect on $\varepsilon^{\infty}$ than the $\chi_{0}$ effect of the chosen density functional. We note that r2SCAN satisfies more exact constraints than PBE does with more accurate total energies in general, suggesting that r2SCAN can have a better $f^{\text{xc}}$ Nazarov and Vignale (2011).
The KS linear response function $\chi_{0}$ describes how the electron density $n$ reacts to small changes in the Kohn-Sham potential $v^{\text{KS}}$. This linear response function is often overestimated in magnitude by SIE, which leads to an overly delocalized and easily perturbable electron density, as well as an underestimated band gap that is inversely proportional to the linear response function.
For CoO, PBE fails to open a band gap, as illustrated in Fig. S1(b). This leads to too large $\chi_{0}$, resulting in dynamical instabilities with the appearance of imaginary frequencies in the phonon calculations as shown in Figure 1(a). In contrast, due to the SIE reduction, r2SCAN predicts an indirect band gap of 0.85 eV, attributed to the splitting between the occupied $t_{2g}$ orbitals and the unoccupied $t_{2g}$ orbitals in the spin-down channel, as shown in Fig. 2(a). This undoubtedly improves $\chi_{0}$ and thus provides an improved phonon dispersion as well as strong EPC effects at the LO and TO phonon modes around the $\Gamma$ point as shown in Figure 1(a). Since a smaller gap generally correlates with a larger $\varepsilon^{\infty}$ Penn (1962); Onishi and Fu (2024), the improved gaps from r2SCAN explain our values of $\varepsilon^{\infty}$ being closer to experimental results as shown in Table. 1.